OpenStudy (anonymous):

Add or subtract the following mixed numbers. First change each mixed number to an equivalent improper fraction. Then find a common denominator, and proceed as before. Leave your answer as an improper fraction. Be sure your answers are reduced to lowest terms. 1 4/9 + 3 1/3 = (options) 53/12 23/12 43/9 23/9

OpenStudy (popitree):

13/9+10/3 =13/9+30/9

OpenStudy (popitree):

proceed like this,, so u have common denominator 9

OpenStudy (texaschic101):

do you know how to change them to improper fractions ?

OpenStudy (anonymous):

can you show me from start to finish?

OpenStudy (texaschic101):

yes...I can 1 4/9.....take your whole number (1) and multiply it by the denominator (9), which equals 9, then add that to your numerator (4), which is 4 + 9 = 13, then put that number over the original denominator (9), and that gives us : 13/9 3 1/3....whole number (3) multiplied by denominator (3), giving us 9, then add it to the numerator (1), giving us 10, and putting it over original denominator (3), giving us : 10/3 do you understand so far ? Because that is how you change a mixed number to an improper fraction. any questions so far ?

OpenStudy (anonymous):

no keep goin

OpenStudy (texaschic101):

so now we have : 13/9 + 10/3....but you cannot add/subtract fractions unless you have a common denominator. So we have to make both fractions have the same denominator. So what number does 9 and 3 both go into evenly ? That would be 9. So we have to make both these fractions have a common denominator of 9. 13/9 -- already has a common denominator of 9, so we do not have to do anything to this one. 10/3 = x/9......we need to find x...3 goes into 9 how many times ? It goes in there 3 times. Now take that 3 and multiply it by the numerator 10, giving us 30...making our fraction 30/9 so now we have : 13/9 + 30/9....now we can add them...just the numerators, the denominators stay the same...giving us 43/9 do you understand ? Because if you don't, I can go over it again.

OpenStudy (texaschic101):

If you do not understand, I can show you another way

OpenStudy (anonymous):

lol yea thanks for the explanation better than school

OpenStudy (texaschic101):

glad I could help :)

OpenStudy (texaschic101):

so you really understand...because, like I said, I can show you another way if needed ?

OpenStudy (anonymous):

well i got this one that is a bit different can you explain? Add or subtract the following mixed numbers using the first method. Be sure your answers are in mixed number format and reduced to lowest terms. 5 5/9 - 3 1/3 = (options) 2 2/9 2 1/3 2 5/6 2 4/9

OpenStudy (texaschic101):

ok...turning to mixed numbers 5 5/9 -- whole number (5) multiplied by denominator (9), which is 45, added to numerator (5), which is 50...put over original denominator (9), which gives us : 50/9 3 1/3 = (3 * 3 + 1) / 3 = 10/3 so now we have : 50/9 - 10/3 --- common denominator is 9 50/9 -- stays the same 10/3 ... here is the other way...we have to find an equivalent fraction for 10/3 that has a denominator of 9. Equivalent fractions can be found by multiplying the numerator and denominator by the same number...for example : 10/3 * 2/2 = 20/6....10/3 is equivalent to 20/6 10/3 * 4/4 = 40/12...10/3 is equivalent to 40/12 we have to find the equivalent fraction for 10/3 that has a denominator of 9. 10 x --- = ---- 3 9 3 times what is 9 ? 3 times 3 is 9...so if we are multiplying the denominato by 3 to get 9, then we have to multiply the numerator by 3 as well. 10 3 30 --- * ---- = ----- 3 3 9 so now we have : 50/9 - 30/9 ... subtract the numerators, do nothing to the denominators. giving us : 20/9 but now we have to turn it back into a mixed number.. 20/9 ... how many times does 9 go into 20 without going over ? That would be 2 times...so 2 is your whole number. 2 * 9 = 18....now take your numerator and subtract 18....20 - 18 = 2....that difference is your numerator....the denominator stays the same. So your mixed number is : 2 2/9....and that is the final answer because it does not reduce. any better ? Or did I just confuse the crap out of you ?

OpenStudy (texaschic101):

I didn't mean to write you a book...I am just trying to explain it to you so you can understand

OpenStudy (texaschic101):

let me know if you have any questions

OpenStudy (anonymous):

it helps alot thanks id give you another medel if i could

OpenStudy (texaschic101):

no worries....I am not here for the medals

OpenStudy (anonymous):

well thanks